OpenStudy (anonymous):

find the derivative of the following function: y=37 arctan (sqrtx).. Please show step by step

OpenStudy (amistre64):

pull the constant; and chain the rest

OpenStudy (anonymous):

as if....

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

RMALIK HAVENT U UNDERSTOOD CHAIN RULE YET? i mean weve done arnd half a dozen qstns for u

OpenStudy (anonymous):

alright

OpenStudy (amistre64):

no need to berate ....

OpenStudy (anonymous):

so do i explain again?

OpenStudy (amistre64):

we explain as many times as it takes :) imo

OpenStudy (anonymous):

and again and again until u finish all the calculus for one lifetime?

OpenStudy (anonymous):

well isnt what this site for? to help

OpenStudy (amistre64):

yes, but its hard at times to tell who is abusing the site; and who is using it

OpenStudy (anonymous):

im not abusing the site i just like to look at steps of the problem

OpenStudy (anonymous):

fyn 1/2(1+x)sqrt (x)

OpenStudy (anonymous):

u treat rootx as x first and then write the derivative

OpenStudy (anonymous):

1/(1 + rootx ^2)

OpenStudy (anonymous):

now multiply it by the derivative of rootx

OpenStudy (amistre64):

tan-1(x) = y means; x = tan(y) ; derive both sides to get 1 = sec^2(y) y' y' = 1/sec^2(y) ; but y = tan-1(x) soo... y' = 1\(sec^2(tan-1(x))) which can be simplified right?

OpenStudy (anonymous):

yes where tan becomes sin over cos

OpenStudy (anonymous):

Personally I like to use a substitution when explaining the chain rule. Since it's often shown as: \[f(g(x))' = f'(x)g'(x)\] So we just need to pick our f and g.

OpenStudy (amistre64):

its been awhile since I tried to step thru it :) Do we translate it from the standard triangle format for it?

OpenStudy (anonymous):

In this case f(a) = 37arctan(a) and g(x) = sqrt(x)

OpenStudy (anonymous):

ok polpak thats good explaination

OpenStudy (anonymous):

I'm sure you can probably find \(g'(x)\) pretty easily by now. So the only question is what's the derivative of arctan(a) ?

OpenStudy (amistre64):

tan(y) = x/1 sec(y) = sqrt(x^2 +1) sec^2(y) = x^2 +1 Dx(tan-1(x)) = 1/(x^2+1) .... right?

OpenStudy (anonymous):

My guess is you have that in a table you're supposed to memorize, but yes amistre64 that's correct.

OpenStudy (amistre64):

yay!! ... that was easier to remember than I thought :) thnx

OpenStudy (anonymous):

Then you just multiply the two derivatives (and simplify if possible)